### v1

@@ 3.2 Astrometry and photometry of protoplanets b and c
The astrometry and contrast of planets and with respect to the star are inferred by using our NEGFC approach (Appendix C), after subtraction of our optimal disk model (Appendix B). These contrast values are then multiplied by the integrated stellar flux values reported in Table 1 to obtain the flux of the protoplanets. We note that the stellar fluxes in the F187N and F480M filters are 25% and 7% brighter than estimated based on the SpeX spectrum presented in Long et al. (2018) and the best-fit SED model to the extracted MIRI-MRS spectrum (Perotti et al. 2023, ), respectively. These differences are compatible with the significant IR variability of the star and inner disk rim (e.g., 25% variation at 5 m between Spitzer and JWST observations; Perotti et al. 2023). These considerations inspire caution regarding protoplanet photometry derived in contrast of the star but based on non-concurrent absolute star photometry measurements.
@@
For each planet, the astrometric values found for the two filters are consistent with each other. The F187N astrometry of planet and candidate is affected by large uncertainties owing to neighbouring disk signals and the overlapping spiral accretion stream. For the F480M images, the large uncertainties also reflect the coarser angular resolution. All estimates are consistent with the expected astrometry of both protoplanets at the date of our observations (given in the last column of Table 2), based on the orbital fits presented in Wang et al. (2021b). Considering our large uncertainties compared to ground-based measurements, we do not attempt new orbital fits in this work.
@@
Figures 2 and 3 show the SED of PDS 70 and , respectively, including our NIRCam data at 1.87 and 4.83 m. The F187N measurement for planet is consistent with predictions from the best-fit atmospheric models obtained with the BT-SETTL, DRIFT-PHOENIX and EXOREM grids (details in Wang et al. 2021b). This does not hold for planet . This discrepancy can have various (non-mutually exclusive) causes, discussed in Appendix F. Here we discuss the hypothesis of significant Pa- line emission from the protoplanet, not captured in the atmospheric model. If the excess signal above the model atmospheric contribution comes only from Pa- emission, the line flux is W m. Given the unknown amplitude of other biases discussed in Appendix F, this estimate is a conservative upper limit. Considering mag (Uyama et al. 2021) and the distance of the system, our constraint on the Pa- luminosity is , where is the solar luminosity. Assuming a planet mass of (Wang et al. 2021b), the models in Aoyama et al. (2021) suggest the mass accretion rate would be yr. The same assumptions for the model and the planet applied to the H flux reported in Haffert et al. (2019) yield an accretion rate yr. If the Aoyama et al. (2021) models are an accurate representation of the accretion process onto giant planets, our results suggest that either additional sources of bias may have a non-negligible effect (e.g., stellar variability or underestimated SPHERE/IFS measurements; Appendix F), or that protoplanets undergo significant accretion variability (e.g. Szulágyi & Ercolano 2020; Casassus & Cárcamo 2022). With a similar reasoning as above, but using the 3- uncertainty on the F187N flux measured for planet and an assumed extinction of (Uyama et al. 2021), we constrain the Pa- luminosity to . Based on the planet accretion models in Aoyama et al. (2021), this translates into an upper mass accretion rate limit of yr for a planet mass (Wang et al. 2021b). This constraint is in line with the accretion rates inferred from the H flux measured for (Haffert et al. 2019), which leads, for the same model and assumptions, to yr. In summary, both measurement uncertainties and additional sources of bias prevent the confirmation of significant Pa- line emission for both protoplanets.
@@
Our new F480M measurement for PDS 70 is consistent with the NACO -band measurement presented in Stolker et al. (2020). Wang et al. (2021b) found that this point was driving the inclusion of an extra blackbody contribution, representative of heated circumplanetary dust, in the atmospheric model with the most support (solid line in Fig. 2). For PDS 70 , our 4.8 m photometry is roughly compatible with the best-fit DRIFT-PHOENIX model, but significantly higher than the best-fit models from the other two grids, which are known to better reproduce the spectra of old L-T dwarfs (e.g., Witte et al. 2011). For both planets, the tentative excess may be attributable to either a warm dusty environment, or ro-vibrational CO line emission from a heated circumplanetary disk (e.g., Oberg et al. 2023). We defer a detailed spectroscopic analysis for a later study including NIRSpec measurements, as the latter have the highest potential to confirm the tentative 4.8 m excesses and constrain their origin.
@@ 3.3 A third protoplanet, a dust clump, or an inner spiral?
The brightest signals in our disk-subtracted F187N image (Fig. 1b) are found near the predicted location of a protoplanet candidate proposed in M19 at the outer edge of the inner disk (referred to as point-like feature therein). The candidate was found in SPHERE/IFS datasets acquired between May 2015 and April 2019. Its reported astrometry, indicated with blue dots in Fig. 1, is consistent with an 13.5 au circular Keplerian orbit in a similar plane as the outer disk, which is assumed for the predicted location shown with dashed circles in Fig. 1. For clarity, we also show the images obtained after also subtracting the estimated flux of protoplanets and (Fig. 1c and f). While inner disk signals are present near the predicted location, there appears to be a significant excess compared to disk signals alone, considering the inner disk emission at a symmetric location with respect to the minor axis to the south-east of the star. While a signal at separation from the star should be considered with caution, the recovery of the inner disk with a geometry close to expected suggests that PSF subtraction residuals are lower than the inner disk signals, hence that the observed excess is likely of circumstellar origin, rather than an artefact. We thus hereafter refer to it as “”, be it tracing a dusty feature or a third protoplanet.
@@
Table 1 reports the astrometry and photometry extracted for in the F187N image. While the F480M image also shows a bright pixel near the expected location, its separation from the star is too small for a reliable contrast estimate. The contrast of derived in the F187N image is about a factor two larger than the median contrast reported in bands in M19. Our estimate is affected by large uncertainties due to the unknown amount of contamination from the inner disk, such that this excess is only marginally significant. Nonetheless, we argue that such excess is expected. The source spectrum is dominated by scattered stellar light at NIR wavelengths (M19), consistent with a very dusty object. In such case, the total intensity scattering phase function should also modulate its brightness along its orbit. Compared to prior epochs, the object is now closer to PA, hence one expects a brighter signal. We estimate a factor 1.6 enhancement in reflected light for between April 2019 and March 2023 based on the variation in outer disk flux measured at PA=PA-PA and PA=PA-PA, where we considered the southwest part of the outer disk for this estimate to avoid any bias from the accretion stream towards the northwest. Within uncertainties, our measured contrast is thus consistent with tracing the same object as M19. This new measurement adds 4 years to the existing 4-year time baseline for orbital coverage, significantly reducing the chance of tracing either a moving illumination effect or the filtered northwestern tip of the inner disk (in M19).
@@
Our independent re-detection of a signal compatible with candidate does not unambiguously confirm its protoplanet nature, but it raises the question of what other physical processes could give rise to a bright and compact NIR signal moving at local Keplerian speed. Given that the spectrum measured in M19 is consistent with tracing scattered stellar light, the authors suggested that it may trace either a transient dust clump or a protoplanet enshrouded in dust. Radiative hydro-dynamical simulations of embedded giant planets suggest they get enshrouded in a dusty circumplanetary disk or envelope which can display a scattered-light dominated spectrum at NIR wavelengths (e.g., Szulágyi et al. 2019). If traces indeed a protoplanet enshrouded in dust, its semi-major axis of 13.5 au would place it near the 1:2:4 mean-motion resonance with planets and ( au and au; Wang et al. 2021b), making especially exciting to follow up. Disentangling the dust clump from the circumplanetary disk or envelope hypothesis will require mid-IR flux measurements (e.g., Chen & Szulágyi 2022). Given the angular separation, this may need the advent of mid-infrared imagers and spectrographs on Extremely Large Telescopes (e.g., ELT/METIS; Brandl et al. 2018). Figure 13 summarizes our proposed interpretation for the main features detected in our NIRCam observations of PDS 70.
@@ A.1 Pre-processing
We first identified remaining bad pixels in the calibrated images through sigma-clipping, using the cube_fix_badpix_clump function of VIP, and corrected them using Gaussian kernel interpolation. We subsequently corrected the imperfect background subtraction performed by the JWST pipeline, which led to negative values, by subtracting the residual background level estimated far from the star. We then found the stellar centroid coordinates using the cube_recenter_dft_upsampling routine in VIP. This routine first leverages the single-step discrete Fourier Transform algorithm presented in Guizar-Sicairos et al. (2008) to find shifts that optimize image cross-correlation throughout the sequence, and then fits a 2D Gaussian model to the mean image of the aligned cube to find the centroid coordinates. Throughout all stages of image processing, all shift and rotation operations were performed in the Fourier plane, the default behavior in VIP111https://github.com/vortex-exoplanet/VIP, as this better preserves pixel intensities (e.g., Larkin et al. 1997).
A.2 PSF subtraction
@@
We investigated different approaches for PSF modeling and subtraction. We first considered pair-wise roll-subtraction (e.g. Schneider et al. 2003) between individual dithered images acquired with each of the two roll angles. We also considered the pair-wise subtraction of the mean image of each of the two sets. We noticed a minor improvement in the residuals near the star which could be due to the spatial undersampling of the images limiting the efficacy of individual pair-wise image subtractions. Both the ‘mean’ and ‘individual’ options are implemented in the roll_sub function of VIP, available as of version 1.6.0. Images are smoothed with a Gaussian kernel the FWHM of which is set to 60% of the observed width of the instrumental PSF. This helps mitigating the undersampling which affects the F187N data and the pixel-to-pixel noise induced by the roll-subtraction approach. In Fig. 4a and e, we show the images obtained with mean roll subtraction on the F187N and F480M data, respectively. Low residuals are achieved owing to the stability of the observed PSF, but significant self-subtraction and geometric distortion of circumstellar signals can also be noted (e.g., by comparison with the IPCA images). Compared to the F480M images shown in Fig. 1, the F480M images in Fig. 4 correspond to the largest common field covered by all dither positions.
@@
In addition to direct roll subtraction, we investigated alternative post-processing methods designed to work in combination with angular differential imaging, namely Principal Component Analysis (PCA; Soummer et al. 2012; Amara & Quanz 2012) and Temporal Reference Analysis for Planets (TRAP; Samland et al. 2021). We used the PCA algorithm implemented in VIP, using as input PCA library for each image, the images obtained at the other roll angle, akin to smartPCA proposed for angular differential imaging (Absil et al. 2013). This yielded very similar results to roll angle subtraction. The TRAP algorithm computes a contrast map. A contrast is computed for each off-axis (RA, DEC)-pair on-sky by simultaneously modeling the pixel lightcurves created by an off-axis PSF at the sky position and the temporal systematics of the pixels affected by this off-axis PSF. For our JWST data there are measurements at two roll angles, such that the off-axis PSF model for each (RA, DEC)-pair results in a step-function with varying jump heights depending on each pixel’s location relative to the off-axis PSF. In practice the contrast map resulting from the forward modeling is a form of local deconvolution with a PSF model that takes into account the field-of-view rotation. This approach is optimized for detecting point sources, but nevertheless works out-of-the-box. However, due to the complex disk structure surrounding the host star it is difficult to locate uncontaminated reference pixels of the host star PSF to model the temporal systematics in the data. For our results, each off-axis PSF pixel lightcurve (step function) is fit simultaneously with only a constant factor. Given the stability of the instrument, this provides good results. However, higher-order noise models or models informed by temporal trends seen in the guide star observations may provide better results in the future. Especially when fitting a forward model that explicitly models the entire scene including disks and planets, overfitting becomes less of a concern. The TRAP analysis is extremely computationally efficient, requiring less than 5 seconds on a laptop. The TRAP images obtained for the F187N and F480M data are shown in Fig. 4b and f, respectively.
@@
As a less aggressive alternative, we performed Reference star Differential Imaging (RDI; e.g., Mawet et al. 2012) using the PCA algorithm implemented in VIP. We tested RDI using observed non-coronagraphic PSFs from the MAST archive as reference library, however this led to strong PSF residuals after subtraction. It is unclear whether this is due to the field dependence of the PSF, undersampling effects, a different spectral slope for the source (for the F480M filter), or a combination of these factors. We therefore also tested creating reference stars using the webbpsf package (v1.1.1; Perrin et al. 2014) as reference PSFs to test RDI. In the latter case, we considered the PSF distortion dependency with location in the field. We built the synthetic PSFs considering a 3972 K black body (Müller et al. 2018) as the stellar spectrum model, anticipating a potential impact of the spectral slope on the PSF for the F480M filter. We also used the optical path differences from the day following the observations, which appear identical to those measured 3 days before the observations. Upon carefully checking the radial profile of the webbpsf synthetic PSFs and the observed PSF of PDS 70, we noticed a broader core for the observed PSF of PDS 70, and residual diffuse hexagonal-shape stellar wings after subtraction. This may suggest a non-negligible contribution from marginally resolved inner disk signals. We also tested PCA-RDI with a library of synthetic PSFs. These were produced separately for each observed dither position on the detector, and considered a range of subpixel shifts around those dither locations (within 0.5px, with steps of 0.05px). Here we attempted to emulate the broader core and undersampling effects. This marginally improved the results, but still yielded strong PSF wing residuals, hence poor-quality final images.
@@
As our RDI attempts were unfruitful, we focused instead on iterative approaches leveraging roll angle diversity. We implemented an iterative roll (IROLL) subtraction algorithm akin to the one presented in Heap et al. (2000), and an Iterative Principal Component Analysis algorithm (IPCA; e.g., Pairet et al. 2021). In either case, circumstellar signals estimated in the processed image obtained at each iteration (e.g., positive residuals above a certain threshold) are removed from the pre-processed images used to create a PSF model (i.e., for each image, the ones obtained at the other roll angle) in the subsequent iteration. IROLL directly uses the images obtained at roll angle as PSF model for images obtained at roll angle (and vice-versa for the roles of and ), and iteratively removes the estimated circumstellar component directly from the roll images. The difference, in the IPCA algorithm, is that for images , principal components are learned from images , and then projected onto the images to produce PSF models for subtraction (and vice-versa). In practice, we used the first principal component with a cube reduced to two images corresponding to the mean image of each roll angle observation. Thus, the only difference between IPCA and IROLL resided in the projection of a normalized mean PSF (i.e., the projected first principal component) in the former case.
@@
While previous IPCA implementations considered all strictly positive residuals (e.g., Pairet et al. 2021; Juillard et al. 2023), here we set an absolute threshold slightly above the noise level achieved in the roll subtraction image, namely 10 MJy/sr and 1 MJy/sr in the F187N and F480M data, respectively. Such absolute threshold should be used with care as it may not be appropriate for all datasets (e.g., images with a strong radial dependency of the residual noise level). Here we noticed it to be relevant given the relatively constant (self-subtracted) noise level in the image. Hence it was particularly efficient at mitigating the propagation of ring-like artefacts (observed in Pairet et al. 2021; Juillard et al. 2023).
@@
The estimated circumstellar signal map at each iteration is smoothed using a thin 2D-Gaussian kernel set to a 1-pixel FWHM before its removal from the roll images used to produce the PSF model images at the subsequent iteration. This intends to capture the spatial correlation of neighbouring pixels, and leads to better recovery of faint signals, originally drowned under the noise level of the roll subtraction image. Our IPCA algorithm iteratively corrects not only for self-subtraction, since circumstellar signals are iteratively removed from the image library used to calculate the principal components, but also for over-subtraction. This is because the model PSF image that is subtracted is built from the projection of principal components onto the original images minus the estimated circumstellar signal map obtained at the previous iteration.
@@
We implemented an automatic convergence criterion based on a user-defined relative tolerance (default 1e-4). If all pixel intensities in the image obtained at the subsequent iteration vary by less than this relative tolerance, the algorithm stops and considers to have converged onto a final image. While this criterion worked for our data, likely helped by the high stability of the PSF, we highlight that there is no mathematical guarantee that this fix-point algorithm will systematically converge in general (see also Juillard et al. 2023). Both IROLL and IPCA converged to a final image within 1000 iterations. While IROLL recovers a significant fraction of the self-subtracted signals, and yields a similar final image as IPCA in the F187N image, this is not the case for the F480M filter image. In the latter case, it does not appear to converge onto an image free of geometric biases (see e.g., Juillard et al. 2022).
@@
As a safety check that the algorithm properly recovers the outer disk, we show in Fig. 5 the Pearson cross-correlation calculated between the F187N image obtained at each iteration of IPCA and our corresponding radiative transfer model of the outer disk, in a mask encompassing the south part of the outer disk (i.e., all the signals to the south of the peak intensity in Fig. 7a; see Sec. B for more details about the disk model). We highlight that this disk model is not used by IPCA, which functions in an agnostic and automated manner. We see that most of the geometric biases are corrected within 200 iterations. Most of the flux is also recovered within a similar number of iterations, with only marginal gains beyond 200 iterations.
@@
The individual and mean roll subtraction, iterative roll subtraction, and iterative PCA algorithms were all implemented in VIP and available as of version 1.6.0. The images obtained with roll subtraction, TRAP, iterative roll subtraction and iterative PCA are shown in Fig. 4.
A.3 Reliability of IPCA

### v2

@@ 3.2 Astrometry and photometry of protoplanets b and c
The astrometry and contrast of planets and with respect to the star were inferred by using our NEGFC approach (Appendix C) after subtraction of our optimal disk model (Appendix B). These contrast values were then multiplied by the integrated stellar flux values reported in Table 1 to obtain the flux of the protoplanets. We note that the stellar fluxes in the F187N and F480M filters are 25% and 7% brighter than estimated based on the SpeX spectrum presented in Long et al. (2018) and the best-fit SED model to the extracted MIRI-MRS spectrum (Perotti et al. 2023), respectively. These differences are compatible with the significant IR variability of the star and inner-disk rim (e.g., 25% variation at 5 m between Spitzer and JWST observations; Perotti et al. 2023). These considerations inspire caution regarding protoplanet photometry derived in contrast of the star that are based on nonconcurrent absolute star photometry measurements.
@@
For each planet, the astrometric values we found for the two filters are consistent with each other. The F187N astrometry of planet and candidate is affected by large uncertainties owing to neighboring disk signals and the overlapping spiral accretion stream. For the F480M images, the large uncertainties also reflect the coarser angular resolution. All estimates are consistent with the expected astrometry of the two protoplanets at the date of our observations (given in the last column of Table 2), based on the orbital fits presented in Wang et al. (2021b). Considering our large uncertainties compared to ground-based measurements, we do not attempt new orbital fits in this work.
@@
Figures 2 and 3 show the SED of PDS 70 and , respectively, including our NIRCam data at 1.87 and 4.83 m. The F187N measurement for planet is consistent with predictions from the best-fit atmospheric models obtained with the BT-Settl, Drift-Phoenix, and EXOREM grids (details in Wang et al. 2021b). This does not hold for planet . This discrepancy can have various (not mutually exclusive) causes that we discuss in Appendix F. Here, we discuss the hypothesis of significant Pa- line emission from the protoplanet that is not captured in the atmospheric model. If the excess signal above the model atmospheric contribution comes from Pa- emission alone, the line flux is W m-2. Given the unknown amplitude of other biases discussed in Appendix F, this estimate is a conservative upper limit. Considering mag (Uyama et al. 2021) and the distance of the system, our constraint on the Pa- luminosity is , where is the solar luminosity. When a planet mass of (Wang et al. 2021b) is assumed, the models in Aoyama et al. (2021) suggest that the mass accretion rate would be yr. The same assumptions for the model and the planet applied to the H flux reported in Haffert et al. (2019) yield an accretion rate yr. If the Aoyama et al. (2021) models are an accurate representation of the accretion process onto giant planets, our results suggest that either additional sources of bias may have a non-negligible effect (e.g., stellar variability or underestimated SPHERE/IFS measurements; Appendix F), or that protoplanets undergo significant accretion variability (e.g. Szulágyi & Ercolano 2020; Casassus & Cárcamo 2022). With a similar reasoning as above, but using the 3 uncertainty on the F187N flux measured for planet and an assumed extinction of (Uyama et al. 2021), we constrain the Pa- luminosity to . Based on the planet accretion models in Aoyama et al. (2021), this translates into an upper mass-accretion rate limit of yr for a planet with a mass (Wang et al. 2021b). This constraint is in line with the accretion rates inferred from the H flux measured for (Haffert et al. 2019), which for the same model and assumptions leads to yr. In summary, the measurement uncertainties and additional sources of bias together prevent us from confirming significant Pa- line emission for the two protoplanets.
@@
Our new F480M measurement for PDS 70 is consistent with the NACO -band measurement presented in Stolker et al. (2020). Wang et al. (2021b) found that this point was driving the inclusion of an additional blackbody contribution that is representative of heated circumplanetary dust in the atmospheric model with the most support (solid line in Fig. 2). For PDS 70 , our 4.8 m photometry is roughly compatible with the best-fit Drift-Phoenix model, but it is significantly higher than the best-fit models from the other two grids, which are known to reproduce the spectra of old L-T dwarfs better (e.g., Witte et al. 2011). For both planets, the tentative excess may be attributable to either a warm dusty environment or to ro-vibrational CO line emission from a heated circumplanetary disk (e.g., Oberg et al. 2023). We defer a detailed spectroscopic analysis to a later study including NIRSpec measurements. The NIRSpec measurements have the highest potential to confirm the tentative 4.8 m excesses and constrain their origin.
@@ 3.3 A third protoplanet, a dust clump, or an inner spiral?
The brightest signals in our disk-subtracted F187N image (Fig. 1b) are found near the predicted location of a protoplanet candidate proposed in M19 at the outer edge of the inner disk (referred to as a point-like feature therein). The candidate was found in SPHERE/IFS datasets acquired between May 2015 and April 2019. Its reported astrometry, indicated with blue dots in Fig. 1, is consistent with an 13.5 au circular Keplerian orbit in a plane similar to that of the outer disk, which is assumed for the predicted location shown with dashed circles in Fig. 1. For clarity, we also show the images that were obtained after the estimated flux of protoplanets and was also subtracted (Fig. 1c and f). While inner-disk signals are present near the predicted location, there appears to be a significant excess compared to the disk signals alone, considering that the inner-disk emission originates in a symmetric location with respect to the minor axis southeast of the star. While a signal at a separation of from the star should be considered with caution, the recovery of the inner disk with a geometry close to expected suggests that the PSF subtraction residuals are lower than the inner-disk signals. This means that the observed excess is likely of circumstellar origin and not an artifact. We therefore refer to it hereafter as “” because it might trace a dusty feature or a third protoplanet.
@@
Table 1 reports the astrometry and photometry we extracted for in the F187N image. While the F480M image also shows a bright pixel near the expected location, its separation from the star is too small for a reliable contrast estimate. The contrast of derived in the F187N image is higher about a factor of 2 than the median contrast reported in bands in M19. Our estimate is affected by large uncertainties due to the unknown amount of contamination from the inner disk, and this excess is therefore only marginally significant. Nonetheless, we argue that this excess is expected. The source spectrum is dominated by scattered stellar light at NIR wavelengths (M19), which is consistent with a very dusty object. In this case, the scattering phase function of the total intensity is also expected to modulate its brightness along its orbit. Compared to prior epochs, the object is now closer to PAβ, and therefore, we would expect a brighter signal. We estimate an enhancement of a factor of 1.6 in reflected light for between April 2019 and March 2023 based on the variation in the flux of the outer disk measured at PA=PAβ-PAd,2019 and PA=PAβ-PAd,2023, where we considered the southwest part of the outer disk for this estimate to avoid any bias from the accretion stream toward the northwest. Within the uncertainties, our measured contrast is therefore consistent with tracing the same object as M19. This new measurement adds 4 years to the existing 4-year time baseline for the orbital coverage, and it significantly reduces the probability that either traces a moving illumination effect or the filtered northwestern tip of the inner disk (in M19).
@@
Our independent redetection of a signal that is compatible with candidate does not unambiguously confirm its protoplanet nature, but it raises the question of which other physical processes might give rise to a bright and compact NIR signal that moves at the local Keplerian speed. The spectrum measured in M19 is consistent with tracing scattered stellar light, and the authors therefore suggested that it might trace either a transient dust clump or a protoplanet enshrouded in dust. Radiative hydrodynamical simulations of embedded giant planets suggest that they become enshrouded in a dusty circumplanetary disk or envelope, which can display a scattered-light dominated spectrum at NIR wavelengths (e.g., Szulágyi et al. 2019). If indeed traces a protoplanet enshrouded in dust, its semimajor axis of 13.5 au would place it near the 1:2:4 mean-motion resonance with planets and ( au and au; Wang et al. 2021b). Follow-up studies of are therefore especially exciting. Distinguishing the hypotheses of a dust clump from the circumplanetary disk or an envelope will require MIR flux measurements (e.g., Chen & Szulágyi 2022). Because of the angular separation, this may need the advent of MIR imagers and spectrographs on extremely large telescopes (e.g., ELT/METIS; Brandl et al. 2018). Figure 13 summarizes our proposed interpretation of the main features detected in our NIRCam observations of PDS 70.
@@ A.1 Preprocessing
We first identified remaining bad pixels in the calibrated images through sigma-clipping using the cube_fix_badpix_clump function of VIP, and we corrected for them using Gaussian kernel interpolation. We subsequently corrected for the imperfect background subtraction performed by the JWST pipeline that led to negative values by subtracting the residual background level estimated far from the star. We then found the stellar centroid coordinates using the cube_recenter_dft_upsampling routine in VIP. This routine first leverages the single-step discrete Fourier transform algorithm presented in Guizar-Sicairos et al. (2008) to find shifts that optimize the image cross-correlation throughout the sequence, and it then fits a 2D Gaussian model to the mean image of the aligned cube to find the centroid coordinates. Throughout all stages of the image processing, all shift and rotation operations were performed in the Fourier plane, the default behavior in VIP111https://github.com/vortex-exoplanet/VIP, as this better preserves pixel intensities (e.g., Larkin et al. 1997).
A.2 Subtraction of the point spread function
@@
We investigated different approaches for the modeling and subtraction of the PSF. We first considered pair-wise roll-subtraction (e.g. Schneider et al. 2003) between individual dithered images acquired with each of the two roll angles. We also considered the pair-wise subtraction of the mean image of each of the two sets. We note a minor improvement in the residuals near the star that might be due to the spatial undersampling of the images which limits the efficacy of individual pair-wise image subtractions. Both the mean and individual options are implemented in the roll_sub function of VIP, which is available as of version 1.6.0. The images were smoothed with a Gaussian kernel, the FWHM of which was set to 60% of the observed width of the instrumental PSF. This helps to mitigate the undersampling that affects the F187N data and the pixel-to-pixel noise induced by the roll-subtraction approach. In Fig. 4a and e, we show the images we obtained with mean roll subtraction on the F187N and F480M data, respectively. Low residuals are achieved owing to the stability of the observed PSF, but significant self-subtraction and geometric distortion of circumstellar signals can also be noted (e.g., by comparison with the IPCA images). Compared to the F480M images shown in Fig. 1, the F480M images in Fig. 4 correspond to the largest common field covered by all dither positions.
@@
In addition to direct roll subtraction, we investigated alternative post-processing methods designed to work in combination with angular differential imaging, namely principal component analysis (PCA; Soummer et al. 2012; Amara & Quanz 2012) and temporal reference analysis for planets (TRAP; Samland et al. 2021). We used the PCA algorithm implemented in VIP and used the images obtained at the other roll angle as input PCA library for each image. This is similar to smartPCA, which was proposed for angular differential imaging (Absil et al. 2013). This yielded very similar results to the roll-angle subtraction. The TRAP algorithm computes a contrast map. A contrast was computed for each off-axis (RA, DEC) pair on-sky by simultaneously modeling the pixel light curves created by an off-axis PSF at the sky position and the temporal systematics of the pixels affected by this off-axis PSF. For our JWST data, there are measurements at two roll angles, such that the off-axis PSF model for each (RA, DEC) pair resulted in a step-function with varying jump heights depending on the location of each pixel relative to the off-axis PSF. In practice, the contrast map resulting from the forward modeling is a form of local deconvolution with a PSF model that takes the field-of-view rotation into account. This approach is optimized for detecting point sources, but it also works out of the box. However, because the disk structure surrounding the host star is complex, it is difficult to locate uncontaminated reference pixels of the host star PSF to model the temporal systematics in the data. For our results, each off-axis PSF pixel light curve (step function) was fit simultaneously with only a constant factor. Because the instrument is stable, this provided good results. However, higher-order noise models or models informed by temporal trends seen in the guide-star observations may provide better results in the future. Especially when fitting a forward model that explicitly models the entire scene including disks and planets, overfitting becomes less of a concern. The TRAP analysis is extremely efficient computationally. It requires less than 5 seconds on a laptop. The TRAP images obtained for the F187N and F480M data are shown in Fig. 4b and f, respectively.
@@
As a less aggressive alternative, we performed reference star differential imaging (RDI; e.g., Mawet et al. 2012) using the PCA algorithm implemented in VIP. We tested RDI using observed noncoronagraphic PSFs from the MAST archive as reference library, but this led to strong PSF residuals after subtraction. It is unclear whether this is due to the field dependence of the PSF, undersampling effects, a different spectral slope for the source (for the F480M filter), or a combination of these factors. We therefore also tested creating reference stars using the webbpsf package (v1.1.1; Perrin et al. 2014) as reference PSFs to test RDI. In the latter case, we considered the PSF distortion dependence on the location in the field. We built the synthetic PSFs considering a 3972 K blackbody (Müller et al. 2018) as the stellar spectrum model, anticipating a potential impact of the spectral slope on the PSF for the F480M filter. We also used the optical path differences from the day following the observations, which appear identical to those measured 3 days before the observations. Upon carefully checking the radial profile of the webbpsf synthetic PSFs and the observed PSF of PDS 70, we noted a broader core for the observed PSF of PDS 70 and residual diffuse hexagonal-shape stellar wings after subtraction. This may suggest a non-negligible contribution from marginally resolved inner-disk signals. We also tested PCA-RDI with a library of synthetic PSFs. These were produced separately for each observed dither position on the detector and considered a range of subpixel shifts around these dither locations (within 0.5px, with steps of 0.05px). Here, we attempted to emulate the broader core and undersampling effects. This marginally improved the results, but still yielded strong PSF wing residuals, and hence, poor-quality final images.
@@
As our RDI attempts were unfruitful, we instead focused on iterative approaches leveraging roll-angle diversity. We implemented an iterative roll (IROLL) subtraction algorithm similar to the one presented in Heap et al. (2000), and an iterative principal component analysis algorithm (IPCA; e.g., Pairet et al. 2021). In either case, circumstellar signals estimated in the processed image obtained at each iteration (e.g., positive residuals above a certain threshold) were removed from the pre-processed images used to create a PSF model (i.e., for each image, those obtained at the other roll angle) in the subsequent iteration. IROLL directly uses the images obtained at roll angle as PSF model for images obtained at roll angle (and vice versa for the roles of and ) and iteratively directly removes the estimated circumstellar component from the roll images. The difference in the IPCA algorithm is that for images , the principal components are learned from images and are then projected onto the images to produce PSF models for subtraction (and vice versa). In practice, we used the first principal component with a cube reduced to two images corresponding to the mean image of each roll-angle observation. Thus, the only difference between IPCA and IROLL resided in the projection of a normalized mean PSF (i.e., the projected first principal component) in the former case.
@@
While previous IPCA implementations considered all strictly positive residuals (e.g., Pairet et al. 2021; Juillard et al. 2023), here, we set an absolute threshold slightly above the noise level achieved in the roll-subtraction image, namely 10 MJy/sr and 1 MJy/sr in the F187N and F480M data, respectively. An absolute threshold like this should be used with care as it may not be appropriate for all datasets (e.g., images with a strong radial dependence on the residual noise level). We realized that this was relevant given the relatively constant (self-subtracted) noise level in the image. It was therefore particularly efficient at mitigating the propagation of ring-like artifacts (observed in Pairet et al. 2021; Juillard et al. 2023).
@@
The estimated circumstellar signal map at each iteration was smoothed using a thin 2D Gaussian kernel set to a one-pixel FWHM before its removal from the roll images that were used to produce the PSF model images at the subsequent iteration. This intends to capture the spatial correlation of neighboring pixels and leads to a better recovery of faint signals that are originally drowned in the noise level of the roll-subtraction image. Our IPCA algorithm iteratively corrects not only for self-subtraction because circumstellar signals are iteratively removed from the image library that is used to calculate the principal components, but also for oversubtraction. This is because the model PSF image that is subtracted is built from the projection of principal components onto the original images minus the estimated circumstellar signal map obtained at the previous iteration.
@@
We implemented an automatic convergence criterion based on a user-defined relative tolerance (default 1e-4). When all pixel intensities in the image obtained at the subsequent iteration vary by less than this relative tolerance, the algorithm stops and considers to have converged onto a final image. While this criterion worked for our data, likely helped by the high stability of the PSF, we highlight that there is no mathematical guarantee that this fix-point algorithm will systematically converge in general (see also Juillard et al. 2023). Both IROLL and IPCA converged to a final image within 1000 iterations. While IROLL recovered a significant fraction of the self-subtracted signals and yielded a similar final image as IPCA in the F187N image, this was not the case for the F480M filter image. In the latter case, it does not appear to converge onto an image free of geometric biases (see e.g., Juillard et al. 2022).
@@
As a safety check that the algorithm properly recovered the outer disk, we show in Fig. 5 the Pearson cross-correlation calculated between the F187N image obtained at each iteration of IPCA and our corresponding radiative transfer model of the outer disk in a mask encompassing the south part of the outer disk (i.e., all the signals south of the peak intensity in Fig. 7a; see Sec. B for more details about the disk model). We highlight that this disk model is not used by IPCA, which functions in an agnostic and automated manner. Most of the geometric biases are corrected within 200 iterations. Most of the flux is also recovered within a similar number of iterations, with only marginal gains beyond 200 iterations.
@@
The individual and mean roll subtraction, iterative roll subtraction, and iterative PCA algorithms were all implemented in VIP and are available as of version 1.6.0. The images obtained with roll subtraction, TRAP, iterative roll subtraction, and iterative PCA are shown in Fig. 4.
A.3 Reliability of the IPCA